# Subtype-Aware Registration of Longitudinal Electronic Health Records

### v1

@@ 3 Simulations
This scenario includes 4 groups of trajectories and each group includes 250 subjects. The trajectories of each group is generated as follows:
$\begin{split}\text{Group 1 (250 subjects):}\quad&Y_{i}^{(1)}(t_{ij})=23+3\sin(0.6(t_{ij}+4))+\epsilon_{ij};\\ \text{Group 1 (250 subjects):}\quad&Y_{i}^{(2)}(t_{ij})=17+3\sin(0.6(t_{ij}+4))+\epsilon_{ij};\\ \text{Group 1 (250 subjects):}\quad&Y_{i}^{(3)}(t_{ij})=20+3\sin(0.6(t_{ij}+4))+\epsilon_{ij};\\ \text{Group 1 (250 subjects):}\quad&Y_{i}^{(4)}(t_{ij})=25-0.5t_{ij}+4\sin(0.9(t_{ij}+3))+\epsilon_{ij}.\end{split}$
In summary, Groups 1, 2, and 3 share a similar shape with different means, while Group 4 has a distinctly different shape but overlaps with other types. This scenario helps evaluate the algorithm’s ability to distinguish between groups with both shape and mean differences, including overlapping patterns.
@@
1 (2 clusters) 69.0% [62.0%, 83.9%] 92.0% [84.9%, 99.5%] 0.40 [0.17, 0.55] 3.13 [1.22, 7.42]
2 (2 clusters) 75.8% [73.1%, 80.0%] 94.5% [90.6%, 98.0%] 0.35 [0.24, 0.49] 1.24 [0.86, 1.93]
3 (3 clusters) 74.0% [71.1%, 78.1%] 91.5% [90.5%, 93.1%] 0.40 [0.34, 0.43] 1.25 [0.87, 1.93]
4 (4 clusters) 69.2% [61.3%, 79.6%] 97.1% [95.7%, 98.5%] 0.37 [0.28, 0.47] 2.63 [1.86, 5.46]
5 (4 clusters) 65.2% [50.7%, 75.9%] 90.6% [87.9%, 93.2%] 0.60 [0.45, 0.77] 4.56 [1.29, 7.57]
6 (4 clusters) 76.0% [69.8%, 80.7%] 93.9% [89.8%, 96.4%] 0.41 [0.30, 0.58] 1.69 [1.14, 2.07]
@@
We use four metrics to evaluate the performance of the algorithm: Accuracy Rate refers to the proportion of subjects for whom the algorithm successfully recovered the correct temporal positions of their longitudinal data; Probability of Error $\leq 1$ Day refers to the proportion of subjects with a result error of less than or equal to 1 day; Mean Absolute Error (MAE) represents the average time difference between the recovered temporal positions and the correct temporal positions for all subjects, and the Runtime indicates the time taken for the algorithm to run once on the simulated dataset. The simulation results, which present the average values and 95% confidence intervals for all four metrics over 500 runs, are shown in Table 1. The visualizations of the simulation results can be found in Figures 1, 3, 4, 5, 6, and 7 in the Appendix.

### v2

@@ A.2 Scenario Design
Three groups with a shared shape and one group with a distinct shape:
$\text{Group 1 (250 subjects):}\quad Y_{i}^{(1)}(t_{ij})$ $=23+3\sin(0.6(t_{ij}+4))+\epsilon_{ij},$
$\text{Group 2 (250 subjects):}\quad Y_{i}^{(2)}(t_{ij})$ $=17+3\sin(0.6(t_{ij}+4))+\epsilon_{ij},$
$\text{Group 3 (250 subjects):}\quad Y_{i}^{(3)}(t_{ij})$ $=20+3\sin(0.6(t_{ij}+4))+\epsilon_{ij},$
$\text{Group 4 (250 subjects):}\quad Y_{i}^{(4)}(t_{ij})$ $=25-0.5t_{ij}+4\sin(0.9(t_{ij}+3))+\epsilon_{ij}.$
Scenario 7:
Heterogeneous progression speeds within Groups 2 and 3, with $s_{i}^{(2a)}=1$ and $s_{i}^{(2b)}=0.7$:
$\text{Group 1 (500 subjects):}\quad Y_{i}^{(1)}(t_{ij})$ $=17+2\sin\!\big(0.6(t_{ij}+4)\big)+\epsilon_{ij},$
$\text{Group 2 (250 subjects):}\quad Y_{i}^{(2a)}\!\big(s_{i}^{(2a)}t_{ij}\big)$ $=2\,t_{ij}+3\sin\!\big(0.6(t_{ij}+4)\big)+\epsilon_{ij},$
$\text{Group 3 (250 subjects):}\quad Y_{i}^{(2b)}\!\big(s_{i}^{(2b)}t_{ij}\big)$ $=2\,t_{ij}+3\sin\!\big(0.6(t_{ij}+4)\big)+\epsilon_{ij}.$
Scenario 8:
Heterogeneous progression speeds across Groups 3 and 4, with $s_{i}^{(3a)}=1$ and $s_{i}^{(3b)}=1.3$:
$\text{Group 1 (300 subjects):}\quad Y_{i}^{(1)}(t_{ij})$ $=20+3\sin\!\big(0.6(t_{ij}+4)\big)+\epsilon_{ij},$
$\text{Group 2 (300 subjects):}\quad Y_{i}^{(2)}(t_{ij})$ $=17+3\sin\!\big(0.6(t_{ij}+4)\big)+\epsilon_{ij},$
$\text{Group 3 (200 subjects):}\quad Y_{i}^{(3a)}\!\big(s_{i}^{(3a)}t_{ij}\big)$ $=16+0.5\,t_{ij}+3\sin\!\big(0.9(t_{ij}+4)\big)+\epsilon_{ij},$
$\text{Group 4 (200 subjects):}\quad Y_{i}^{(3b)}\!\big(s_{i}^{(3b)}t_{ij}\big)$ $=16+0.5\,t_{ij}+3\sin\!\big(0.9(t_{ij}+4)\big)+\epsilon_{ij}.$
An illustrative example for Scenario 1 is shown in Figure 4. Additional visualizations for all scenarios are provided in Figures 511.
A.3 Results
We use four metrics to evaluate the performance of the algorithm: Accuracy Rate refers to the proportion of subjects for whom the algorithm successfully recovered the correct temporal positions of their longitudinal data; Probability of Error $\leq 1$ Day refers to the proportion of subjects with a result error of less than or equal to 1 day; Mean Absolute Error (MAE) represents the average time difference between the recovered temporal positions and the correct temporal positions for all subjects, and the Runtime indicates the time taken for the algorithm to run once on the simulated dataset. The simulation results, which present the average values and 95% confidence intervals for all four metrics over 500 runs, are shown in Table 8. The visualizations of the simulation results can be found in Figure 4 and Figures 5, 6, 7, 8, 9, 10, and 11 in the supplementary materials.
@@ A.3 Results
1 (2 clusters) 65.2% [61.7%, 74.2%] 83.4% [76.5%, 97.6%] 0.56 [0.29, 0.69] 0.77 [0.61, 1.11]
2 (2 clusters) 75.5% [71.8%, 79.1%] 93.5% [89.7%, 96.4%] 0.37 [0.27, 0.47] 0.72 [0.57, 1.19]
3 (3 clusters) 78.1% [76.3%, 81.4%] 91.2% [90.2%, 92.8%] 0.36 [0.31, 0.39] 0.69 [0.61, 0.81]
4 (4 clusters) 73.4% [70.7%, 82.9%] 96.9% [95.6%, 98.2%] 0.33 [0.25, 0.38] 0.71 [0.56, 1.19]
5 (4 clusters) 70.8% [60.3%, 79.4%] 90.9% [87.9%, 93.5%] 0.54 [0.39, 0.68] 1.13 [0.62, 1.94]
6 (4 clusters) 79.7% [74.7%, 83.6%] 93.9% [89.8%, 96.4%] 0.37 [0.27, 0.53] 0.72 [0.57, 1.19]
7 (3 clusters) 72.4% [69.4%, 76.1%] 92.7% [91.7%, 94.5%] 0.43 [0.37, 0.48] 0.73 [0.58, 1.21]
8 (3 clusters) 71.9% [60.9%, 75.9%] 86.8% [79.7%, 89.3%] 0.50 [0.42, 0.71] 0.74 [0.64, 0.93]
@@
Figure 5: Simulation 2 trajectories
Figure 6: Simulation 3 trajectories
Figure 7: Simulation 4 trajectories
Figure 8: Simulation 5 trajectories
Figure 9: Simulation 6 trajectories
Figure 10: Simulation 7 trajectories
Figure 11: Simulation 8 trajectories